OpenStudy (anonymous):

A jar of mayonnaise is shaped like a cylinder. The radius of the jar is 1.5 in. and its height is 7 in. The mayonnaise is on sale for $0.08 per in3. What is the cost of the jar of mayonnaise? Use 3.14 to approximate pi. $

OpenStudy (anonymous):

Plz help will give fan Medal !!

OpenStudy (thebadyboy):

what do u think

OpenStudy (anonymous):

Idk Lol

OpenStudy (igreen):

First find the volume of the mayonnaise jar. Volume of a Cylinder: \(V = \pi r^2 h\) \(V = (3.14)(1.5^2)(7)\) Simplify that

OpenStudy (anonymous):



OpenStudy (thebadyboy):



OpenStudy (igreen):

Then we can divide by 0.08

OpenStudy (anonymous):

yes

OpenStudy (thebadyboy):

he is correct

OpenStudy (anonymous):

So by what

OpenStudy (igreen):

@MAli13chineta It's a cylinder, not a circle.

OpenStudy (thebadyboy):

yah we know that

OpenStudy (igreen):

Multiply 3.14 * 1.5 * 1.5 * 7 @TLHernandez

OpenStudy (thebadyboy):

but thx

OpenStudy (anonymous):

49.455

OpenStudy (anonymous):

So whats the amswer

OpenStudy (anonymous):

???

OpenStudy (igreen):

Sorry, multiply 49.455 * 0.08

OpenStudy (igreen):

Can you do that? @TLHernandez that will be your answer.

OpenStudy (anonymous):

yes

OpenStudy (igreen):

Okay, tell me what you get.

OpenStudy (anonymous):

3.9564

OpenStudy (igreen):

Yep, so the jar is approximately $3.96

OpenStudy (anonymous):

thanks coould you help me with a few more ?

OpenStudy (igreen):

Sure, just close this one and open a new one.

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.